The length of a rectangle is 7 meters longer than its width. What is the width of this rectangle if its perimeter is equal to 130 meters?
Answer by moocow(3)

easy question
width = x
length = x+7
x + x + x+7 + x+7 = 130
4x+17=130
4x = 113
x = 29
plug it in
width = 29
length = 29+7 = 36!
CHECK 36 + 36 + 29 + 29 = 130